Add the following unsigned binary numbers:

01110101 + 00111011

Subtract the following binary numbers using 2's Complement

01110101 - 00111011

SHOW STEPS
Answer #1
1)
01110101
00111011
---------
10110000
---------
Answer: 10110000

2)
01110101 - 00111011
-00111011
follow below steps to calculate - of 00111011
1)  Flip all the bits(0's to 1, 1's to 0)
    00111011 is flipped to 11000100
2)  Add 1 to the above result
    11000100 + 1 = 11000101
so, 01110101 - 00111011 = 01110101 + 11000101
   01110101
   11000101
   ---------
(1)00111010
   ---------
Answer: 00111010 with a carry of 1